### v1

# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

Abstract.
The streaming bipartite graph is extensively used to model the dynamic relationship between two types of entities in many real-world applications, such as movie recommendations, location-based services, and online shopping. Since it contains abundant information, discovering the dense subgraph with high structural cohesiveness (i.e., community detection) in the bipartite streaming graph is becoming a valuable problem. Inspired by this, in this paper, we study the structure of community on the butterfly motif in the bipartite graph. We propose a novel problem, named Community Detection over Streaming Bipartite Network (CD-SBN), which aims to retrieve qualified communities with user-specific query keywords and high structural cohesiveness at snapshot and continuous scenarios. In particular, we formulate the user relationship score in the weighted bipartite network via the butterfly pattern and define a novel $(k,r,\sigma)$-bitruss as the community structure. To efficiently tackle the CD-SBN problem, we design effective pruning strategies to rule out false alarms of $(k,r,\sigma)$-bitruss and propose a hierarchical synopsis to facilitate the CD-SBN processing. Due to the dynamic of streaming bipartite networks, we devise an efficient procedure for incremental graph maintenance. We develop an efficient algorithm to answer the snapshot and continuous CD-SBN query by traversing the synopsis and applying the pruning strategies. With extensive experiments, we demonstrate the efficiency and effectiveness of our proposed CD-SBN processing approach over real/synthetic streaming bipartite networks.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/L1ANLab/CD-SBN.
(a) $G_{t-1}$ at timestamp $(t-1)$
(b) $G_{t}$ at timestamp $t$
Figure 1. An example of community detection in streaming click bipartite networks $G_{t}$.
1. Introduction
Nowadays, the management of bipartite graphs that involve two distinct types of nodes (e.g., customers and products, visitors and check-in locations, users and clicked websites) has been extensively studied in many real applications such as Online Recommendation from user-product transaction graphs (Fan et al., 2022; Weng et al., 2022; Kou et al., 2023), location-based services with visitor-location check-in graphs (Hang et al., 2018; Zhou et al., 2021a; Wang et al., 2017; Kim et al., 2022), and behavior analysis in user-Webpage click graphs (Sun et al., 2009; Neria et al., 2017).
The community detection (CD) over such bipartite graphs has recently become increasingly important for bipartite graph analysis and mining (Barber, 2007; Bouguessa and Nouri, 2021; Yen and Larremore, 2020; Zhou and Amini, 2019), which uncovers hidden and structurally cohesive bipartite subgraphs (or groups). The CD problem over bipartite graphs has many useful real applications, such as online product recommendation and marketing, user behavior analysis, malicious user group identification, and so on.
Although many prior works (Wang et al., 2023; Abidi et al., 2023a; Wang et al., 2021; Zhang et al., 2024) studied the CD problem in static bipartite graphs, real-world bipartite graphs are usually dynamically changing over time (e.g., updates of purchase transactions, check-ins, or website browsing). Therefore, in this paper, we will consider the Community Detection over Streaming Bipartite Network (CD-SBN), upon graph updates (e.g., edge insertions/deletions or edge weight changes).
Below, we give a motivating example of our CD-SBN problem in the application of detecting malicious user groups (communities) over a user-website click bipartite graph.
Example 0.
(Anomaly Detection for the Cybersecurity) Figure 1(a) illustrates an example of a clickstream bipartite network $G_{t-1}$ at timestamp $(t-1)$, which contains 3 user vertices, $u_{1}\sim u_{3}$, 2 website vertices, $v_{1}$ and $v_{2}$, and 5 edges $(u_{i},v_{a})$ between two types of nodes, user $u_{i}$ and website $v_{a}$ (for $1\leq i\leq 3$ and $1\leq a\leq 2$). Here, each website vertex $v_{a}$ (for $a=1,2$) has a set of keywords that represent the website’s features (e.g., {bank, finance} for a bank website $v_{1}$). Moreover, each edge $(u_{i},v_{a})$ (for $i=1\sim 3$ and $a=1,2$) is associated with an integer weight, indicating the frequency that user $u_{i}$ accessed website $v_{a}$ for the past five minutes.
As shown in Figure 1(a), to detect malicious groups of users accessing sensitive/suspicious websites, the network security officer may want to specify some query keywords of websites (e.g., “bank” and “malware”), and identify some group (community), $g_{1}$, of users (e.g., $u_{2}$ and $u_{3}$) who frequently visit some common websites with query keywords (e.g., bank website $v_{1}$ and hacking website $v_{2}$). The officer will warn those suspicious users in the resulting community $g_{1}$ and/or take immediate actions (e.g., recording the evidence for the police).
Figure 1(b) shows the streaming bipartite network $G_{t}$ at timestamp $t$, with a newly inserted edge $(u_{1},v_{1})$ and an expired edge $(u_{3},v_{2})$ (i.e., user $u_{3}$ has not visited website $v_{2}$ for the past five minutes). In this case, at timestamp $t$, the community in $G_{t}$ is changed to $g_{2}$ (rather than $g_{1}$ at timestamp $(t-1)$), as circled by the orange dashed line. $\blacksquare$
In addition to the example above, the CD-SBN problem has many other real applications. For example, in the customer-product transaction bipartite graph, we can use the CD-SBN results to find communities of customers who recently have a purchasing behaviors for online advertising and marketing. Similarly, in the visitor-location check-in bipartite graph (e.g., from Yelp (Inc., 2024)), we can identify a group of visitors who frequently check in at some common points of interest (POIs) and provide them with group buying coupons/discounts (e.g., Groupon (Groupon, 2024)).
In this paper, we formulate and tackle the CD-SBN problem, which obtains all bipartite communities with the user-specified query keywords (e.g., POIs’ features) and high structural cohesiveness (e.g., small distance between any two users and high structural score of the community). In particular, we formally define the community semantics in the context of bipartite graphs, that is, a keyword-aware and structurally dense bipartite subgraph (i.e., a so-called $(k,r,\sigma)$-bitruss that contains query keywords, as will be described in Section 2.4). We consider both snapshot and continuous scenarios of our CD-SBN problem, which detect communities over a snapshot of streaming bipartite network $G_{t}$ at timestamp $t$, or continuously monitor the CD-SBN answer set for each registered community constraint upon streaming graph updates, respectively.
Due to the large scale of bipartite graphs and rapid streaming graph changes, efficient processing of the snapshot and continuous CD-SBN is rather challenging. In order to tackle the challenges, in this paper, we propose effective pruning strategies with respect to the community constraints (e.g., query keywords, community radius, edge supports, and community scores) to significantly reduce the CD-SBN problem space. We design a hierarchical synopsis to effectively facilitate the candidate community search, and develop efficient snapshot and continuous algorithms to retrieve or incrementally maintain actual community answers (via our proposed synopsis and pruning methods), respectively.
In this paper, we make the following major contributions.
  1. (1)
    We formally define the problem of the community detection over streaming bipartite network (CD-SBN) containing snapshot and continuous queries in Section 2.
  2. (2)
    We present a framework to efficiently process the two types of queries with incremental updates of the bipartite streaming graph in Section 3.
  3. (3)
    We design effective community-level pruning strategies to reduce the search space of the CD-SBN problem in Section 4.
  4. (4)
    We propose a novel hierarchical synopsis to facilitate CD-SBN query processing and devise an efficient procedure for incremental graph maintenance in Section 5.
  5. (5)
    We develop an efficient algorithm with effective synopsis-level pruning strategies to answer the snapshot CD-SBN query and another algorithm to maintain the result set of a continuous CD-SBN query with low computational cost in Section 6.
  6. (6)
    We demonstrate the efficiency and effectiveness of our CD-SBN processing approach through extensive experiments over synthetic/real-world graphs in Section 7.
Section 8 reviews the related work on community search/ detection on unipartite graphs and static/streaming bipartite graphs. Finally, Section 9 concludes this paper.
2. Problem Definition
Section 2.1 provides the data model for streaming bipartite networks. Section 2.2 defines the basic units that indicate the relationship between two users. Section 2.3 introduces the relationship score and the $(k,r,\sigma)$-bitruss to measure the cohesiveness of a subgraph. Section 2.4 formally defines our problem of the community detection over streaming bipartite networks (CD-SBN).
2.1. Streaming Bipartite Networks
Bipartite Graphs: We first define an undirected and weighted bipartite graph as follows.
Definition 0.
(Bipartite Graph, $G$) A bipartite graph, $G$, is represented by a quadruple $(U(G),L(G),E(G),\Phi(G))$, where $U(G)$ and $L(G)$ are two disjoint sets (types) of vertices, $E(G)$ is a set of edges, $e_{u,v}$, between vertices $u\in U(G)$ and $v\in L(G)$, and $\Phi(G)$ is a mapping function: $U(G)\times L(G)\rightarrow E(G)$.
Each vertex $v\in L(G)$ has a set, $v.K$, of keywords, and each edge $e_{u,v}\in E(G)$ is associated with a weight $w_{u,v}$.
Bipartite graphs have been widely used in many real applications, such as movie recommendations for users (Grujić, 2008), recommending places of interest (POIs) to tourists (Lang et al., 2022), or online product shopping recommendation for customers (Chen et al., 2022; Shan and Zhao, 2017; Yoon et al., 2020).
For simplicity, this paper considers vertices $u$ in the upper vertex layer $U(G)$ as users in real-world applications, and vertices $v$ in the lower vertex layer $L(G)$ as items that are associated with keyword sets $v.K$ such as movie types, POI features, and product descriptions. Moreover, each edge $e_{u,v}$ from user $u$ to item $v$ is associated with a weight $w_{u,v}$, indicating the user-item interaction frequency such as ##\## of movie views for movie recommendations, ##\## of POI visits by users, ##\## of product purchases for online shopping, and so on.
Streaming Bipartite Networks: Below, we introduce the bipartite network in the streaming scenario.
Definition 0.
(Streaming Bipartite Network, $G_{t}$) A streaming bipartite network $G_{t}$ consists of an initial bipartite graph $G_{0}=(U(G_{0}),L(G_{0}),E(G_{0}),\Phi(G_{0}))$, and an ordered sequence of update items $S=\{p_{1},p_{2},\ldots,p_{t},\ldots\}$, where each item $p_{t}=(e_{u,v},t)$ indicates an update operator of weight $w_{u,v}$ for an inserted edge $e_{u,v}$ arriving at timestamp $t$. A sliding window, $W_{t}$, of size $s$ in a streaming bipartite network $G_{t}$ contains the most recent $s$ update items $p_{t-s+1}$, $p_{t-s+2}$, $\ldots$, and $p_{t}$, over bipartite graph $G_{t}$, where $t$ is the current timestamp.
In Definition 2.2, we consider a sliding window, $W_{t}$, of $s$ update items over a streaming bipartite network $G_{t}$. Each update item can be the insertion of a new edge/interaction between users and items (or equivalently the increase of the edge weight $w_{u,v}$).
After applying update items in sliding window $W_{t}$ to the bipartite graph $G$, we can obtain the latest snapshot, $G_{t}$, of the streaming bipartite network at timestamp $t$.
Sliding Window Maintenance of Streaming Bipartite Networks: We maintain the sliding window $W_{t}$ upon dynamic updates in $G_{t}$. When a new item $p_{t+1}$ arrives at timestamp $(t+1)$, we will add $p_{t+1}$ to $W_{t}$ and remove the expired item $p_{t-s+1}$ from $W_{t}$, which yields a new sliding window $W_{t+1}$.
Insertion of a New Item $p_{t+1}$. For the insertion of a new item $p_{t+1}$ $=(e_{u,v},t+1)$, we consider two cases:
  • When both vertices $u$ and $v$ exist in $G_{t}$ (i.e., $u\in U(G_{t})$ and $v\in L(G_{t})$), we add edge $e_{u,v}$ to $G_{t}$ and increment edge weight $w_{u,v}$ by 1 (i.e., $w_{u,v}=w_{u,v}+1$), and;
  • When either of vertices $u$ and $v$ does not exist in $G_{t}$ (i.e., $u\notin U(G_{t})$ and/or $v\notin L(G_{t})$), we add new vertex(es) $u$ and/or $v$ to $U(G_{t})$ and/or $L(G_{t})$, respectively, insert a new edge $e_{u,v}$ into $G_{t}$, and set edge weight $w_{u,v}$ to 1 (i.e., $w_{u,v}=1$).
Expiration of an Old Item $p_{t-s+1}$. When an old item $p_{t-s+1}=(e_{u,v},t-s+1)$ expires, we decrement the edge weight $w_{u,v}$ by 1 (i.e., $w_{u,v}=w_{u,v}-1$). If $w_{u,v}=0$ holds, we remove the edge $e_{u,v}$ from $G_{t}$ (keeping two ending vertices $u$ and $v$).
2.2. Basic Patterns in the Bipartite Graph
Figure 2. An example of basic patterns over bipartite network.
In this subsection, we give the definitions of basic patterns, wedge and butterfly (Gou and Zou, 2023), in the bipartite graph $G$.
Wedge: A wedge is formally defined as follows:
Definition 0.
(Wedge, $\angle(u_{i},v,u_{j})$) Given two user vertices $u_{i},u_{j}$ $\in U(G)$ and an item vertex $v\in L(G)$ in a bipartite graph $G$, a wedge (Gou and Zou, 2023), $\angle(u_{i},v,u_{j})$, is a path $u_{i}\rightarrow v\rightarrow u_{j}$, where $u_{i}$, $v$, and $u_{j}$ are called start-, middle-, and end-vertex of the wedge, respectively. The weight of a wedge $\angle(u_{i},v,u_{j})$ is given by:
(1) $w_{\angle(u_{i},v,u_{j})}=\min\{w_{u_{i},v},w_{u_{j},v}\}.$
Based on Definition 2.3, a wedge $\angle(u_{i},v,u_{j})$ in the bipartite graph can be two audiences watching the same movie, two travelers visiting the same site, or two customers purchasing the same product.
The weight of the wedge $\angle(u_{i},v,u_{j})$ takes the minimum edge weights between $w_{u_{i},v}$ and $w_{u_{j},v}$, indicating the intensity (or frequency) of maximally common interest/behavior $v$ (e.g., movie, POI site, or product) between two users $u_{i}$ and $u_{j}$.
In the example of Figure 2, $\angle(u_{1},v_{1},u_{2})$ is one of the wedges, with weight $w_{\angle(u_{1},v_{1},u_{2})}=\min\{2,4\}=2$.
Butterfly: Two wedges form a butterfly, as defined below:
Definition 0.
(Butterfly (Gou and Zou, 2023), $\Join(u_{i},u_{j},v_{a},v_{b})$) Given two user vertices $u_{i},u_{j}\in U(G)$ and two item vertices $v_{a},v_{b}\in L(G)$ in a bipartite graph $G$, a butterfly, $\Join(u_{i},u_{j},v_{a},v_{b})$, is a complete bipartite subgraph containing wedges $\angle(u_{i},v_{a},u_{j})$ and $\angle(u_{i},v_{b},u_{j})$.
The butterfly structure (as given in Definition 2.4) has been widely used in many algorithms and applications (Sanei-Mehri et al., 2018; Xu et al., 2022b; Zhou et al., 2021b) to detect communities with high cohesiveness in bipartite graphs. In Figure 2, $\Join(u_{2},u_{3},v_{2},v_{3})$ is an example of a butterfly.
The Score of a Butterfly: The butterfly is a basic graphlet pattern of bipartite graphs. Below, we define a butterfly score to measure the relationship among user and item vertices in the butterfly.
Definition 0.
(The Butterfly Score) The score, $w_{\Join(u_{i},u_{j},v_{a},v_{b})}$, of a butterfly $\Join(u_{i},u_{j},v_{a},v_{b})$ is given by:
$\vspace{-3ex}w_{\Join(u_{i},u_{j},v_{a},v_{b})}$ $=$ $w_{\angle(u_{i},v_{a},u_{j})}\cdot w_{\angle(u_{i},v_{b},u_{j})}$
$=$ $\min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}\cdot\min\{w_{u_{i},v_{b}},w_{u_{j},v_{b}}\}.$
In Definition 2.5, we consider two wedges $\angle(u_{i},v_{a},u_{j})$ and $\angle(u_{i},v_{b},$ $u_{j})$ in the butterfly $\Join(u_{i},u_{j},v_{a},v_{b})$, and define the score of the butterfly by multiplying their corresponding weights, that is, $\min\{w_{u_{i},v_{a}},$ $w_{u_{j},v_{a}}\}\cdot\min\{w_{u_{i},v_{b}},$ $w_{u_{j},v_{b}}\}$. Intuitively, we treat the weight of each wedge as the frequency of duplicate wedges, and the score of the butterfly is given by the count of duplicate butterflies (i.e., the multiplication of these two wedge weights).
For the previous example in Figure 2, the score, $w_{\Join(u_{2},u_{3},v_{2},v_{3})}$, of the butterfly $\Join(u_{2},u_{3},v_{2},v_{3})$ can be calculated by: $\min\{5,1\}\cdot\min\{6,3\}=1\times 3=3$.
2.3. The $(k,r,\sigma)$-Bitruss Community in a Bipartite Graph
In this subsection, we first define the user relationship score between any two user vertices $u_{1}$ and $u_{2}$ in a bipartite graph $G$ and then formalize the $(k,r,\sigma)$-bitruss community with high structural cohesiveness in $G$.
The Relationship Score of a User Vertex Pair: Based on the score of a butterfly (as given in Eq. (2.5)), we can define the user relationship score of a user vertex pair $(u_{1},u_{2})$ as follows.
Definition 0.
(User Relationship Score, $score_{u_{i},u_{j}}$) Given two vertices $u_{i},u_{j}\in U(G_{t})$ in a bipartite graph $G_{t}$, the user relationship score, $score_{u_{i},u_{j}}$, between $u_{i}$ and $u_{j}$ in $G_{t}$ can be calculated by:
(3) $score_{u_{i},u_{j}}(G_{t})=\sum_{\forall v_{a},v_{b}\in L(G_{t})}{w_{\Join(u_{i},u_{j},v_{a},v_{b})}}.$
In Definition 2.6, the user relationship score $score_{u_{i},u_{j}}$ between two users $u_{i}$ and $u_{j}$ is defined as the summed butterfly score for all butterflies containing vertices $u_{i}$ and $u_{j}$. A higher relationship score indicates a stronger relationship between users $u_{i}$ and $u_{j}$.
[…]
The $\mathbf{(k,r,\sigma)}$-Bitruss: We next define a $(k,r,\sigma)$-bitruss community in the bipartite graph $G$ with high structural and user relationship cohesiveness.
Definition 0.
($(k,r,\sigma)$-Bitruss) Given a bipartite graph $G_{t}=(U(G_{t}),L(G_{t}),E(G_{t}),\Phi(G_{t}))$, a center vertex $u_{c}\in U(G_{t})$, a threshold $k$ of butterfly number, a maximum radius $r$, and a threshold, $\sigma$, of the user relationship score, a $(k,r,\sigma)$-bitruss is a connected subgraph, $g$, of $G_{t}$ (denoted as $g\subseteq G_{t}$), such that:
  • (Support) for each edge $e_{u,v}\in E(g)$, the edge support $sup(e_{u,v})$ (defined as the number of butterflies containing edge $e_{u,v}$) is not smaller than $k$;
  • (Radius) for any user vertex $u_{i}\in U(g)$, we have $dist(u_{c},u_{i})\leq 2r$, and;
  • (Score) for any two user vertices $u_{i},u_{j}\in\angle(u_{i},v,u_{j})$, we have $score_{u_{i},u_{j}}(g)\geq\sigma$,
where $dist(u_{c},u_{i})$ is the shortest path distance between $u_{c}$ and $u_{i}$ in bipartite subgraph $g$.
2.4. The Problem of the Community Detection Over Streaming Bipartite Network (CD-SBN)
The CD-SBN Problem: We are now ready to define the problem of detection for communities over a streaming bipartite network upon dynamic updates.
Definition 0.
(Community Detection Over Streaming Bipartite Network, CD-SBN) Given a streaming bipartite network $G_{t}$, a support threshold $k$, a maximum radius $r$, a threshold, $\sigma$, of the user relationship score, and a set, $Q$, of query keywords, the problem of the community detection over streaming bipartite network (CD-SBN) retrieves a result set, $R$, of subgraphs $g_{i}$ of $G_{t}$ (i.e., $g_{i}\subseteq G_{t}$), such that:
  • (Keyword Relevance) for any item vertex $v_{i}\in L(g_{i})$, its keyword set $v_{i}.K$ must contain at least one query keyword in $Q$ (i.e., $v_{i}.K\cap Q\neq\emptyset$), and;
  • (Structural and User Relationship Cohesiveness) bipartite subgraph $g_{i}$ is a $(k,r,\sigma)$-bitruss (as given in Definition 2.7).
Note that, the CD-SBN problem over streaming bipartite network in Definition 2.8 has two scenarios, one-time snapshot and continuous CD-SBN versions. In particular, the one-time snapshot CD-SBN problem finds CD-SBN communities over a snapshot, $G_{t}$, of a streaming bipartite graph at timestamp $t$, whereas the continuous CD-SBN continuously monitors CD-SBN communities upon edge updates (i.e., $S$) in $G_{t}$, for some registered community constraints in Definition 2.8. The former version can be used to obtain initial CD-SBN community answers when CD-SBN predicates are online specified, and the latter one dynamically maintains CD-SBN community results upon graph updates.
Challenges: The processing of snapshot/continuous CD-SBN queries is quite challenging. One straightforward method to tackle this problem is to enumerate all community candidates and check whether or not these candidates satisfy the CD-SBN constraints. However, this method is quite inefficient. First, the checking of CD-SBN community predicates is very costly and involves the constraints of keywords, graph structural cohesiveness, and user relationship scores. Second, there are many possible CD-SBN community candidates, which incurs high computation costs to enumerate and refine. Third, upon graph updates over a sliding window $W_{t}$, it is also challenging to incrementally maintain the dynamically changing graph $G_{t}$ and efficiently update the CD-SBN query answers. Table 1 shows the commonly used symbols and their descriptions in this paper.
3. The CD-SBN Processing Framework
Figure 1 illustrates a CD-SBN processing framework for answering snapshot/continuous CD-SBN queries, which consists of three components, initialization, graph incremental maintenance, and CD-SBN query processing.
To support the CD-SBN stream processing, the initialization component prepares some auxiliary data over initial bipartite graph $G_{0}$ at timestamp 00, including a keyword bit vector $v.BV$ for hashing keywords in item vertices, wedge-score-related data $X_{u_{i},u_{l}}$ and $Y_{u_{i},u_{l}}$ for the butterfly score calculation, support upper bound $ub\_sup(e_{u,v})$, and synopses, $Syn$ (lines 1-4).
At each timestamp $t$, the sliding window $W_{t}$ includes a new item $p_{t}$ (i.e., the edge insertion) and removes an expired item $p_{t-s}$ (i.e., the edge deletion). Therefore, the graph incremental maintenance component will compute an incremental update of wedge-score-related data, $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$ (given by Eqs. (5.2) and  (5.2), respectively), and obtain the updated wedge-score-related data $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ at timestamp $t$ (lines 5-6). Similarly, for those affected edges $e^{\prime}$ (due to edge insertion/deletion), we will also update support upper bounds $ub\_sup(e^{\prime})$ (line 7) and synopsis $Syn$ (line 8).
For the CD-SBN query processing component, the snapshot CD-SBN algorithm traverses the synopsis $Syn$ to retrieve candidate communities $g$, by applying our proposed pruning strategies, and refines the resulting candidate communities $g$ to obtain/return actual CD-SBN answers in $R$ (lines 9-12). Moreover, the continuous CD-SBN algorithm incrementally maintains CD-SBN community answers in $R$ by removing those communities (with the expired edge) and inserting new CD-SBN communities (due to the edge insertion) (lines 13-15).
Table 1. Symbols and descriptions.
Symbol Description
$G$ a bipartite graph
$G_{t}$ a streaming bipartite network at timestamp $t$
$U(G)$ a set of user vertices $u$ in the upper layer of bipartite graph $G$
$L(G)$ a set of item vertices $v$ in the lower layer of bipartite graph $G$
$E(G)$ a set of edges $e_{u,v}$ for $u\in U(G)$ and $v\in L(G)$
$S$ an ordered sequence of update items $p_{t}=(e_{u,v},t)$
$u$ (or $u_{i}$, $u_{j}$) a user vertex in $U(G)$ of bipartite graph $G$
$v$ (or $v_{a}$, $v_{b}$) an item vertex in $L(G)$ of bipartite graph $G$
$W_{t}$ a sliding window of size $s$ at timestamp $t$
$w_{u,v}$ the edge weight of edge $e_{u,v}$
$\angle(u_{i},v,u_{j})$ a wedge with $u_{i},u_{j}\in U(G)$ and $v\in L(G)$
$\Join(u_{i},u_{j},v_{a},v_{b})$ a butterfly with $u_{i},u_{j}\in U(G)$ and $v_{a},v_{b}\in L(G)$
$w_{\Join(u_{i},u_{j},v_{a},v_{b})}$ the score of a butterfly $\Join(u_{i},u_{j},v_{a},v_{b})$
$k$ the support threshold for a $(k,r,\sigma)$-bitruss
$sup(e)$ the number of butterflies containing edge $e$
$r$ the maximum radius for a $(k,r,\sigma)$-bitruss
$\sigma$ the relationship score threshold for a $(k,r,\sigma)$-bitruss
$v_{i}.K$ the keyword set for each vertex $v_{i}\in L(G)$
Input: i) a streaming bipartite network $G_{t}$; ii) a set, $Q$, of query keywords; iii) a support threshold $k$; iv) a maximum radius $r$; v) a threshold, $\sigma$, of user relationship score
Output: a set, $R$, of CD-SBN communities
1
// Initialization ($t=0$)
2 hash keywords in $v.K$ into a keyword bit vector $v.BV$ for each item vertex $v\in L(G_{0})$
3 calculate the wedge-score-related data $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ (as given in Eq. (4.5)) for each pair of user vertices $u_{i},u_{j}\in U(G_{0})$
4 compute the support upper bound $ub\_sup(e_{u,v})$ for each edge $e_{u,v}\in E(G_{0})$
5 construct a hierarchical synopsis, $Syn$, over $G_{0}$
6
7for each timestamp $t$ ($>0$) do
       // Graph Incremental Maintenance
8       compute $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$ (as given in Eqs. (5.2) and  (5.2), resp.) and obtain the updated $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$
9       update support upper bounds $ub\_sup(e^{\prime})$ for relevant edges $e^{\prime}$ (upon edge insertion of $p_{t}$ and deletion of $p_{t-s}$)
10       update synopsis $Syn$
11
      // CD-SBN Query Processing
12       for each snapshot CD-SBN query do
13             traverse the synopsis $Syn$ by applying keyword, support, and score upper bound pruning strategies to retrieve candidate communities $g$
14             refine candidate communities $g$ and obtain actual CD-SBN results $R$
15             return $R$
16
17
18      for each continuous CD-SBN query do
19             remove the expired communities from $R$
20            insert new CD-SBN communities into $R$
21
22
Algorithm 1 The CD-SBN Query Process Framework
4. Pruning Strategies
This section details the community-level pruning strategies (see Section 3) to reduce the search space of the CD-SBN problem.
4.1. Keyword Pruning
Based on the CD-SBN problem definition (as given Definition 2.8), any item vertex $v_{i}$ in the lower layer of a candidate bipartite subgraph $g$ should contain at least one query keyword in $Q$. In the sequel, we present an effective keyword pruning, which filters out item vertices $v_{i}$ in subgraph $g$ that do not contain any query keywords in $Q$.
Lemma 0.
(Keyword Pruning) Given a query keyword set $Q$ and a candidate bipartite subgraph $g$, any item vertex $v_{i}\in L(g)$ can be safely pruned from the subgraph $g$, if $v_{i}.K\cap Q=\emptyset$ holds, where $v_{i}.K$ is the keyword set associated with item vertex $v$.
Proof.
If $v_{i}.K\cap Q=\emptyset$ holds for an item vertex $v_{i}\in L(g)$, then vertex $v_{i}$ does not contain any query keyword in $Q$, which violates the keyword relevance constraint (as given in Definition 2.8). Hence, the item vertex $v_{i}$ cannot be in the CD-SBN subgraph answer and can thus be safely pruned, which completes the proof. ∎
4.2. Support Pruning
As mentioned in Definition 2.8, since the CD-SBN community should be a $(k,r,\sigma)$-bitruss, the support, $sup(e_{u,v})$, of edge $e_{u,v}$ should not be smaller than $k$ (i.e., $sup(e_{u,v})\geq k$). Assume that we can obtain an upper bound, $ub\_sup(e_{u,v})$, of support $sup(e_{u,v})$ for an edge $e_{u,v}\in E(g)$. We provide the following support pruning lemma to rule out edges in a candidate bipartite subgraph $g$ with low supports.
Lemma 0.
(Support Pruning) Given a support threshold $k$ and a candidate bipartite subgraph $g$, the edge $e_{u,v}\in E(g)$ can be safely pruned from subgraph $g$ if it satisfies $ub\_sup(e_{u,v})<k$.
Proof.
According to Definition 2.7, the support, $sup(e_{u,v})$, is defined as the number of butterflies containing a specific edge. Thus, each edge $e$ must be reinforced by at least $k$ butterflies in a $(k,r,\sigma)$-bitruss. Since $ub\_sup(e_{u,v})$ is an upper bound of support $sup(e_{u,v})$, we have $sup(e_{u,v})\leq ub\_sup(e_{u,v})$. Moreover, from the assumption of the lemma that $ub\_sup(e_{u,v})<k$ holds, by the inequality transition, we have $sup(e_{u,v})\leq ub\_sup(e_{u,v})<k$. Hence, the support $sup(e_{u,v})$ of the edge $e_{u,v}$ is below the threshold $k$, and can be safely pruned, which completes the proof. ∎
Discussions on How to Obtain the Upper Bound of Support $\mathbf{ub\_sup(e_{u,v})}$. In the offline phase, the upper bound of support needs to be computed to enable the support pruning. Since part of item vertices in data graph $G_{t}$ will be filtered out by the keyword pruning, where the rest of the vertices induce a subgraph $g$, the number of butterflies containing the edge $e_{u,v}$ in $g$ is smaller than or equal to that in $G_{t}$. Thus, the support of $e_{u,v}$ in $G_{t}$ can be considered as a supper upper bound $ub\_sup(e_{u,v})$ in subgraph $g$ defined below.
(4) $\vspace{-1.5ex}ub\_sup(e_{u,v})=\sum_{\forall v_{x}\in N(u)-\{v\}}{|N(v)\cap N(v_{x})-\{u\}|},\vspace{-0.5ex}$
where $N(v)$ is the neighbor vertices set of vertex $v$. With $u\in U(G_{t})$ and $v\in L(g)$, the number of butterflies containing $e_{u,v}$ in $G_{t}$ can be calculated as the sum of the number of common neighbors for $v$ and each neighbor $v_{x}$ of $u$.
4.3. Layer Size Pruning
Since the support of a bipartite subgraph is closely related to the subgraph size, we can utilize filter out a candidate community $g$ with small sizes of upper/lower layers (i.e., $|U(g)|$ and $|L(g)|$), with respect to the support threshold $k$.
Lemma 0.
(Layer Size Pruning) Given a support threshold $k$ and a candidate bipartite subgraph $g=(U(g),L(g),E(g),\Phi(g))$, subgraph $g$ can be safely pruned, if it satisfies the condition that $(|U(g)|-1)\cdot(|L(g)|-1)<k$.
Proof.
According to Definition 2.7, the support of each edge $e$ in $(k,r,\sigma)-bitruss$ must be higher than the support threshold $k$, which means that each edge $e$ must be contained by at least $k$ butterflies. For each edge, $e_{u_{i},v_{a}}$, any pair of other user vertex and item vertex, $(u_{x},v_{y})$, can form a butterfly, $\Join(u_{i},u_{x},v_{a},v_{y})$, with the two ending vertices of edge $e_{u_{i},v_{a}}$. Assume that $g$ is a fully connected bipartite graph, the edge $e_{u_{i},v_{a}}$ is supported by pairs of vertices $|U-\{u_{i}\}|\times|L-\{v_{a}\}|$ so that its support is $sup(e_{u_{i},v_{a}})=(|U(g)|-1)\cdot(|L(g)|-1)$. However, not every pair of vertices is connected to $u_{i}$ or $v_{a}$ in general, hence it holds that $sup(e_{u_{i},v_{a}})\leq(|U(g)|-1)\cdot(|L(g)|-1)$. If $(|U(g)|-1)\cdot(|L(g)|-1)<k$ holds, we have $sup(e_{u_{i},v_{a}})\leq(|U(g)|-1)\cdot(|L(g)|-1)<k$. Thus, edge $e_{u_{i},v_{a}}$ can be safely pruned, which completes the proof. ∎
4.4. Radius Pruning
In the definition of $(k,r,\sigma)$-bitruss, the radius parameter $r$ ensures that the distances between the center vertex and other user vertices are smaller than or equal to $2r$. Thus, we devise a radius pruning strategy to discard a user vertex $u_{i}$ whose distance, $dist(u_{i},u_{q})$, to the center vertex $u_{q}$ is greater than $2r$.
Lemma 0.
(Radius Pruning) Given a candidate bipartite subgraph $g$ (centered at user vertex $u_{q}$) and a radius $r$, a user vertex $u_{i}\in U(g)$ can be safely pruned from subgraph $g$, if it holds $dist(u_{i},u_{q})>2r$.
Proof.
For the candidate subgraph $g$ centered at the user vertex $u_{q}$, if $dist(u_{i},u_{q})>2r$ holds for a user vertex $u_{i}\in U(g)$, then $u_{i}$ violates the radius constraint. Therefore, we should prune the user vertex $u_{i}$ from $g$, which completes the proof. ∎
4.5. Score Upper Bound Pruning
Based on Definition 2.7, any pair of user vertices belonging to a wedge in a $(k,r,\sigma)$-bitruss must have the user relationship score greater than or equal to $\sigma$, where $\sigma$ is a score threshold specified by users online. In the sequel, we propose a score upper bound pruning strategy to eliminate the user vertex(es) with low scores.
Lemma 0.
(Score Upper Bound Pruning) Given a community $g$ and a candidate user vertex $u_{i}\in U(g)$, the user vertex $u_{i}$ can be safely pruned, if there exists a user vertex $u_{j}\in U(g)$ and a wedge $\angle(u_{i},v,u_{j})\subseteq g$, such that $ub\_score_{u_{i},u_{j}}(g)<\sigma$ holds, where $ub\_score_{u_{i},u_{j}}(g)$ is an upper bound of score $score_{u_{i},u_{j}}(g)$.
Proof.
Given a candidate user vertex $u_{i}$ and a user vertex $u_{j}\in U(g)$, since $ub\_score_{u_{i},u_{j}}(g)$ is a score upper bound, it holds that $score_{u_{i},u_{j}}(g)\leq ub\_score_{u_{i},u_{j}}(g)$. From the assumption of the lemma, if $ub\_score_{u_{i},u_{j}}(g)<\sigma$ holds, by the inequality transition, we have $score_{u_{i},u_{j}}(g)\leq ub\_score_{u_{i},u_{j}}(g)<\sigma$, violating the score constraint of the $(k,r,\sigma)$-bitruss (given in Definition 2.7). Thus, the candidate user vertex $u_{i}$ can be safely pruned, which completes the proof. ∎
Discussions on How to Obtain the Score Upper Bound,
$ub\_score_{u_{i},u_{j}}(g)$:
To enable the score upper bound pruning, we need to calculate the score upper bound, $ub\_score_{u_{i},u_{j}}(g)$, for the user pair $(u_{i},u_{j})$ in subgraph $g$ that is connected to some common-neighbor item vertices with the desired (query) keywords. While Eq. (3) computes the user relationship score $score_{u_{i},u_{j}}(G_{t})$, by considering all possible common-neighbor item vertices $v\in L(G_{t})$ of users $u_{i}$ and $u_{j}$, and ignoring the constraint of query keywords (i.e., $Q$), the score $score_{u_{i},u_{j}}(G_{t})$ is essentially an upper bound of $score_{u_{i},u_{j}}(g)$ (since $g\subseteq G$ holds). Therefore, we can use $score_{u_{i},u_{j}}(G_{t})$ as the score upper bound $ub\_score_{u_{i},u_{j}}(g)$ (i.e., $ub\_score_{u_{i},u_{j}}(g)=score_{u_{i},u_{j}}(G_{t})$).
Computation of the Score $score_{u_{i},u_{j}}(G_{t})$: Now the only remaining issue is how to efficiently compute the user relationship score, $score_{u_{i},u_{j}}(G_{t})$, which is given by the summed score for all butterflies containing $u_{i}$ and $u_{j}$ by combining Eqs. (3) and  (2.5).
Based on the multinomial theorem (Bolton, 1968) with power $n=2$, we have:
$\left(\sum_{i=1}^{m}{o_{i}}\right)^{2}$ $=$ $\sum_{i=1}^{m}{o_{i}}^{2}+2\sum_{1\leq i<j\leq m}{o_{i}o_{j}},$
which can be rewritten as:
(5) $\sum_{1\leq i<j\leq m}{o_{i}o_{j}}$ $=$ $\frac{1}{2}\left(\left(\sum_{i=1}^{m}{o_{i}}\right)^{2}-\sum_{i=1}^{m}{o_{i}}^{2}\right).$
Based on Eq. (5), the user relationship score, $score_{u_{i},u_{j}}(G_{t})$, between $u_{i}$ and $u_{j}$ (as given in Eq. (3)) can be rewritten as:
$score_{u_{i},u_{j}}(G_{t})=\sum_{\forall a<b,v_{a},v_{b}\in L(G_{t})}{w_{\Join(u_{i},u_{j},v_{a},v_{b})}}$
$=$ $\sum_{\forall a<b,v_{a},v_{b}\in L(G_{t})}{w_{\angle(u_{i},v_{a},u_{j})}\cdot{w_{\angle(u_{i},v_{b},u_{j})}}}$
$=$ $\frac{1}{2}\left(\left(\sum_{\forall v\in L(G_{t})}{w_{\angle(u_{i},v,u_{j})}}\right)^{2}-\sum_{\forall v\in L(G_{t})}\left(w_{\angle(u_{i},v,u_{j})}\right)^{2}\right)$
$=$ $\frac{1}{2}\left(\left(\sum_{\forall v\in L(G_{t})}{\min\{w_{u_{i},v},w_{u_{j},v}\}}\right)^{2}-\sum_{\forall v\in L(G_{t})}\left(\min\{w_{u_{i},v},w_{u_{j},v}\}\right)^{2}\right)$
$=$ $\frac{1}{2}\left(X_{u_{i},u_{j}}^{2}-Y_{u_{i},u_{j}}\right).$
From Eq. (4.5), in order to calculate the user relationship score $score_{u_{i},u_{j}}(G_{t})$ for each user pair $u_{i}$ and $u_{j}$, we only need to maintain two terms, $X_{u_{i},u_{j}}=\sum_{\forall v\in L(G_{t})}{\min\{w_{u_{i},v},w_{u_{j},v}\}}$ and $Y_{u_{i},u_{j}}=\sum_{\forall v\in L(G_{t})}\left(\min\{w_{u_{i},v},w_{u_{j},v}\}\right)^{2}$. Upon graph updates in stream $S$, we can incrementally and efficiently update $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ online (as will be discussed later in Section 5.2).
5. Auxiliary Data Initialization and Incremental Maintenance
This section details the initialization stage (lines 1-4 of Algorithm 1) that computes auxiliary data over an initial bipartite graph $G_{0}$ in Section 5.1, and constructs a synopsis, $Syn$, for such auxiliary data in Section 5.3, as well as incremental maintenance of auxiliary data and synopsis (lines 5-8 of Algorithm 1) in Sections 5.2 and  5.4, respectively.
5.1. Auxiliary Data Initialization
To facilitate efficient online community detection, we will compute some auxiliary data over the initial graph $G_{0}$ in the initialization phase, which can enable our proposed pruning strategies (as mentioned in Section 4).
  • (Item Keyword Bit Vector) a keyword bit vector, $v.BV$, of size $B$ for each item vertex $v\in L(G_{0})$, whose elements $v.BV[i]$ are $1$ if a keyword in $v.K$ is hashed to the $i$-th position (otherwise, $v.BV[i]=0$);
  • (User Keyword Bit Vector) a keyword bit vector, $u.BV$, of each user vertex $u$, which is a bit-OR of bit vectors $v.BV$ for all $u$’s neighbor item vertices $v\in N(u)$ (i.e., $u.BV=\bigvee_{\forall v\in N(u))}v.BV$);
  • (Support Upper Bound) an upper bound, $ub\_sup(e_{u,v})$ (as given by Eq. (4)), of the support $sup(e_{u,v})$, for each edge $e_{u,v}\in E(G_{0})$, and;
  • (Aggregated Wedge Scores) the summed wedge score, $X_{u_{i},u_{j}}$ $=\sum_{\forall v\in L(G_{t})}$ $w_{\angle(u_{i},v,u_{j})}$, and the summation over the squares of wedge scores, $Y_{u_{i},u_{j}}=$ $\sum_{\forall v\in L(G_{t})}(w_{\angle(u_{i},v,u_{j})})^{2}$, for each user pair $(u_{i},u_{j})\in U(G_{0})\times U(G_{0})$ and item vertices $v\in N(u_{i})\cap N(u_{j})$.
5.2. Incremental Maintenance of Auxiliary Data
In this subsection, we illustrate how to incrementally maintain auxiliary data of graph $G_{t}$ for each user vertex, upon changes of the update items in the sliding window $W_{t}$. As mentioned in Section 2.1, the sliding window $W_{t}$ at a new timestamp $t$ (from $W_{t-1}$) needs to insert a new item $p_{t}$ and remove an expired old item $p_{t-s}$. Upon such updates, we need to incrementally compute the edge weights in the graph $G_{t}$, as well as the auxiliary data in the synopsis $Syn$.
Maintenance of Keyword Bit Vectors: When an edge $e_{u_{i},v_{a}}$ is inserted or deleted in the bipartite graph $G_{t}$, we need to update the keyword bit vector $u_{i}.BV$. Specifically, for the insertion of a new edge $e_{u_{i},v_{a}}$, the new keyword bit vector is given by $u_{i}.BV\bigvee v_{a}.BV$; for the deletion of an expired edge $e_{u_{i},v_{a}}$, we recompute the keyword bit vector $u_{i}.BV=\bigvee_{\forall v\in N(u_{i})}v.BV$.
Maintenance of Support Upper Bounds: At timestamp $t$, due to the insertion of $p_{t}$ or the deletion of $p_{t-s}$, an edge $e_{u_{i},v_{a}}$ can be added or deleted, respectively. Consequently, we need to update the support upper bounds of the edges in all butterflies $\Join(u_{i},u_{j},v_{a},v_{b})$ containing the edge $e_{u_{i},v_{a}}$. When inserting a new edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we increase the support upper bound $ub\_sup(e)$ by 1. Similarly, when deleting an edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we decrease the support upper bound $ub\_sup(e)$ by 1.
Maintenance of User Relationship Scores: Since the user relationship score $score_{u_{i},u_{j}}(G_{t})$ is calculated using $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ (as given by Eq. (4.5)), we maintain these two terms incrementally. Specifically, we compute the differences, $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$, of $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$, respectively, between timestamps $(t-1)$ and $t$. Upon updating the edge $e_{u_{i},v_{a}}$ at timestamp $t$, the edge weight $w_{u_{i},v_{a}}$ has an increase/decrease of 1. Accordingly, the wedge weight $w_{\angle(u_{i},v_{a},u_{j})}$ ($=min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$) is updated with $w_{\angle(u_{i},v_{a},u_{j})}$ $+$ $\lambda_{\angle(u_{i},v_{a},u_{j})}$, where $\lambda_{\angle(u_{i},v_{a},u_{j})}$$\in$$\{-1,0,1\}$ (depending on the change of $min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$).
Here, there are three possible cases for the $\lambda_{\angle(u_{i},v_{a},u_{j})}$ value:
  • (Case 1) if the weight $w_{u_{i},v_{a}}$ of edge $e_{u_{i},v_{a}}$ is increased by 1 and $w_{u_{i},v_{a}}+1\leq w_{u_{j},v_{a}}$, then we have $\lambda_{\angle(u_{i},v_{a},u_{j})}=1$;
  • (Case 2) if the weight $w_{u_{i},v_{a}}$ of edge $e_{u_{i},v_{a}}$ is decreased by 1 and $w_{u_{i},v_{a}}-1<w_{u_{j},v_{a}}$, then we have $\lambda_{\angle(u_{i},v_{a},u_{j})}=-1$;
  • (Case 3) for the remaining case, we have $\lambda_{\angle(u_{i},v_{a},u_{j})}=0$.
The Computation of Terms $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$: We provide the formulae of the two terms $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$ below.
$\Delta X_{u_{i},u_{j}}$
$=$ $X_{u_{i},u_{j}}^{(t)}-X_{u_{i},u_{j}}^{(t-1)}$
$=$ $\left(\sum_{\forall v\in L(G),v\neq v_{a}}{w_{\angle(u_{i},v,u_{j})}}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})\right)-\sum_{\forall v\in L(G)}{w_{\angle(u_{i},v,u_{j})}}$
$=$ $\lambda_{\angle(u_{i},v_{a},u_{j})},$
where $X_{u_{i},u_{j}}^{(t)}$ and $X_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $X_{u_{i},u_{j}}$ at timestamps $t$ and $(t-1)$, respectively.
From Eq. (5.2), for any user pair $(u_{i},u_{j})$, the increase/decrease of $X_{u_{i},u_{j}}$ is equal to that of $w_{\angle(u_{i},v_{a},u_{j})}$ (i.e., $\lambda_{\angle(u_{i},v_{a},u_{j})}$).
$\Delta Y_{u_{i},u_{j}}$
$=$ $Y_{u_{i},u_{j}}^{(t)}-Y_{u_{i},u_{j}}^{(t-1)}$
$=$ $\left(\sum_{\forall v\in L(G),v\neq v_{a}}\left(w_{\angle(u_{i},v,u_{j})}\right)^{2}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})^{2}\right)$
$-\sum_{\forall v\in L(G)}\left(w_{\angle(u_{i},v,u_{j})}\right)^{2}$
$=$ $2\lambda_{\angle(u_{i},v_{a},u_{j})}w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})}^{2},$
where $Y_{u_{i},u_{j}}^{(t)}$ and $Y_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $Y_{u_{i},u_{j}}$ at timestamps $t$ and $(t-1)$, respectively.
As given in Eq. (5.2), the increment/decrement of $Y_{u_{i},u_{j}}$ can be computed via $w_{\angle(u_{i},v_{a},u_{j})}$ and $\lambda_{\angle(u_{i},v_{a},u_{j})}$.
Time Complexity Analysis: When a new edge $e_{u_{i},v_{a}}$ is inserted, the updated keyword bit vector is given by $u_{i}.BV\bigvee v_{a}.BV$, which takes $O(1)$ cost; when an edge $e_{u_{i},v_{a}}$ is deleted, we can re-compute the keyword bit vector $u_{i}.BV=\bigvee_{\forall v\in N(u_{i})}v.BV$, which has the time complexity $O(|N(u_{i})|)$.
For an updated edge $e_{u_{i},v_{a}}$, we update the support upper bound of edges connected to $v_{a}$ and $v_{b}$, where $v_{b}\in N(u_{i})\cap N(u_{j})-\{v_{a}\},u_{j}\in N(v_{a})$. Thus, the worst-case time cost is $O(|N(v_{a})|\cdot|N(u_{i})|)$.
Upon the update of edge $e_{u_{i},v_{a}}$, for any user pair $(u_{i},u_{j})$, the time complexity of updating $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ is given by $O(1)$. The worst-case time complexity is given by $O(|N(v_{a})|)$, where $N(v_{a})$ is a set of user neighbors $u_{j}$ of the item vertex $v_{a}$.
Therefore, the total time complexity of the graph increment maintenance is given by $O(|N(v_{a})|\cdot|N(u_{i})|)$.
5.3. The Synopsis Construction
In this subsection, we present a hierarchical tree synopsis, $Syn$, over auxiliary data for the dynamic graph $G_{t}$ (i.e., initial graph $G_{0}$ or its subsequent version at timestamp $t$) to accelerate online CD-SBN query processing.
The Data Structure of Synopsis, $Syn$: We design a hierarchical synopsis, $Syn$, over the initial bipartite graph $G_{0}$, where each synopsis node, $M$, has multiple entries $M_{i}$, and each of them corresponds to a partition of $G_{0}$. In detail, the hierarchical synopsis consists of two types of nodes: leaf and non-leaf nodes.
Leaf Nodes: Each leaf node $M$ contains entries, each in the form of a user vertex $u\in U(G_{t})$, associated with auxiliary data, $u.agg$. For each possible radius $r\in[1,r_{max}]$, we calculate and store aggregates in the form $(u.BV_{r},u.ub\_sup_{r},u.ub\_score_{r})$ as follows:
  • a keyword bit vector, $u.BV_{r}$, of user vertex $u$, which is a bit-OR of bit vectors $u_{i}.BV$ for all user vertices $u_{i}$ in $(2r)$-hop subgraph of $u$ (i.e., $u.BV_{r}=\bigvee_{\forall u_{i}\in hop(u,2r))}u_{i}.BV$);
  • a support upper bound, $u.ub\_sup_{r}$ of support upper bounds for edges in the $(2r)$-hop subgraph of $u$ (i.e.,$u.ub\_sup_{r}=max_{\forall e\in hop(u,2r)}{ub\_sup(e)}$), and;
  • a user relationship score upper bound, $u.ub\_score_{r}$, of scores, $score_{u_{i},u_{j}}(G_{t})$, for all user pairs $(u_{i},u_{j})$ within the $(2r)$-hop subgraph of user $u$ (i.e., $u.ub\_score_{r}=max_{\forall u_{i},u_{j}\in hop(u,2r)}\\ score_{u_{i},u_{j}}(G_{t})$).
Non-Leaf Nodes: Each non-leaf node $M$ in the synopsis $Syn$ contains entries $M_{i}$, each associated with the following aggregated data $M_{i}.agg$ (w.r.t. each possible radius $r\in[1,r_{max}]$).
  • an aggregated keyword bit vector $M_{i}.BV_{r}=\bigvee_{\forall u\in M_{i}}u.BV_{r}$;
  • the maximum support upper bound $M_{i}.ub\_sup_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_sup_{r}}$, and;
  • the maximum score upper bound $M_{i}.ub\_score_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_score_{r}}$.
Bottom-up Construction of the Synopsis $\mathbf{Syn}$: We construct the hierarchical synopsis, $Syn$, as follows. For the initial bipartite graph $G_{0}$ (i.e., timestamp $t=0$), we start from each vertex $u_{i}\in U(G_{0})$ and perform a breadth-first search (BFS) to extract a $(2r)$-hop subgraph, $hop(u_{i},2r)$, centered at $u_{i}$ with radius $2r$ (for $r\in[1,r_{max}]$). Next, for each subgraph $hop(u_{i},2r)$, we compute its auxiliary data, $u_{i}.agg$. We sort the vertices in $G_{0}$ according to their summed score upper bounds for all possible radii (i.e., $\sum_{\forall r\in[1,r_{max}]}{u.ub\_score_{r}}$), and divide them into multiple partitions of the same size $\gamma$ (i.e., vertices with similar score upper bounds are grouped together), which are leaf nodes of the synopsis $Syn$ (as discussed in Section 5.3). Then, in a bottom-up manner, we recursively obtain non-leaf nodes $M$ of the synopsis $Syn$, by grouping leaf or non-leaf nodes $M_{i}$ (of the same size) on the lower level. The score upper bound associated with node $M$ is computed by aggregating the score upper bounds of its child nodes (i.e., $\sum_{\forall r\in[1,r_{max}]}$ $M_{i}.ub\_score_{r}$). The recursive synopsis construction terminates until one root node, $root(Syn)$, is obtained. After building the tree structure of the synopsis, we will recursively calculate aggregates, $M_{i}.agg$, of entries $M_{i}$ in leaf/non-leaf nodes of synopsis $Syn$. In addition, we keep an inverted list, $inv\_list$, for user vertices $u_{i}\in U(G_{t})$, where each user vertex $u_{i}$ is associated with an ordered list of node IDs from root to a leaf node in synopsis $Syn$ in which $u_{i}$ resides. This inverted list, $inv\_list$, can be used for accessing a particular user vertex for the synopsis maintenance (as discussed in Section 5.4).
5.4. Incremental Maintenance of the Synopsis
In this subsection, we discuss how to identify the affected vertices and nodes in the synopsis $Syn$ and update their aggregates in $Syn$.
Identification of the Affected Vertices and Synopsis Nodes: When an item from $S$ arrives or expires in $W_{t}$ (i.e., an edge $e_{u_{i},v_{a}}$ changes its weight $w_{u_{i},v_{a}}$ or is inserted/removed into/from graph $G_{t-1}$), the auxiliary data, $u.agg$, of some user vertices need to be updated. Therefore, our first goal is to identify those potentially affected vertices (with aggregate updates), which are user vertices in the $(2r)$-hop subgraph of $u_{i}$ for all possible radii $r\in[1,r_{max}]$.
Specifically, if edge $e_{u_{i},v_{a}}$ is inserted or deleted, then we consider those users $u\in hop(u_{i},2r)$ within $(2r)$ hops away from $u_{i}$ as the affected vertices, which need to update keyword bit vectors $BV_{r}$. Moreover, since the support of each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$ may increase/decrease, user vertices, $u$, within $(2r)$-hop from $u_{i}$ may be affected (as $(2r)$-hop subgraph of $u$ may contain those edges $e$). Furthermore, the user relationship score between $u_{i}$ and each vertex $u_{j}\in N(v_{a})$ may change. Thus, we need to update the scores for those user vertices $u_{j}$ whose $(2r)$-hop subgraphs contain both $u_{i}$ and $u_{j}$.
Finally, in synopsis $Syn$, we identify those (non-)leaf nodes that contain the affected user vertices, whose aggregated data will be updated.
Maintenance of Leaf and Non-leaf Nodes: After identifying the affected vertices/nodes, we update their aggregates for all possible radii $r\in[1,r_{max}]$. Assume that an edge $e_{u_{i},v_{e}}$ is inserted or expired, and $u$ is an affected vertex. For the edge insertion, we merge $u_{i}.BV$ into $u.BV_{r}$, update $u.ub\_sup_{r}$ and $u.ub\_score_{r}$ for each affected vertex $u$, and update aggregates of the affected synopsis nodes $M$ containing $u$ (via $inv\_list$). For the edge expiration, we re-aggregate the data, $u.BV_{r}$, $u.ub\_sup_{r}$, and $u.ub\_score_{r}$, from $(2r)$-hop of each affected vertex $u$, and we re-aggregate the data in the affected nodes containing $u$ (if the re-aggregated data on the lower-level affected nodes have changed).
Maintenance of the Newly Added User Vertices: When an edge is inserted with a new user vertex $u_{new}$, we first compute the aggregates, $u_{new}.agg$ for this new user $u_{new}$ (within $(2r)$ hops from $u_{new}$) and add vertex $u_{new}$ to a leaf node $M_{i}$ in synopsis $Syn$ (containing a user vertex with the closest score upper bound). Next, we update the aggregate data of non-leaf nodes containing leaf node $M_{i}$.
[…]
6. CD-SBN query Processing
This section presents pruning strategies over synopsis $Syn$ to filter out unqualified synopsis nodes and provide algorithms for snapshot and continuous CD-SBN processing by accessing the synopsis.
6.1. Pruning via Synopsis
Keyword Pruning for Synopsis Entries: We first discuss how to use the aggregated keyword bit vector, $M_{i}.BV$, to rule out a synopsis entry $M_{i}$, if no query keyword in $Q$ exists in the vertices under $M_{i}$.
Lemma 0.
(Synopsis Keyword Pruning) Given a synopsis entry $M_{i}$ and a set, $Q$, of query keywords, entry $M_{i}$ can be safely pruned, if it holds $M_{i}.BV\bigwedge Q.BV=\boldsymbol{0}$, where $Q.BV$ is a bit vector hashed from query keywords in $Q$.
Proof.
Since $M_{i}.BV$ is an aggregated keyword bit vector computed by $\bigvee_{\forall u\in M_{i}}u.BV$, the $f(w)$-th bit position of $M_{i}.BV$ with the value 00 means that the keyword $w$ is not contained in the keyword set, $v.K$, of any item vertex $v$, where $v\in N(u)$ and the $f(\cdot)$ is a hash function. Therefore, if $M_{i}.BV\wedge Q.BV=\boldsymbol{0}$ holds, no query keywords in $Q$ belong to $M_{i}$, which violates the keyword constraint in Definition 2.8 (i.e., $v.K\cap Q\neq\emptyset$). In this case, the synopsis entry $M_{i}$ cannot contain any vertices in the CD-SBN query answer and thus can be safely pruned, which completes the proof. ∎
Support Pruning for Synopsis Entries: Next, we utilize the maximum support upper bound, $M_{i}.ub\_sup_{r}$, and the query support threshold $k$ to filter out a synopsis entry $M_{i}$.
Lemma 0.
(Synopsis Support Pruning) Given a synopsis entry $M_{i}$ and a support threshold $k$, entry $M_{i}$ can be safely pruned, if it holds that $M_{i}.ub\_sup_{r}<k$, where $M_{i}.ub\_sup_{r}$ is the maximum edge support upper bound in all $(2r)$-hop subgraphs of vertices under $M_{i}$.
Proof.
Since $M_{i}.ub\_sup_{r}$ is the maximum edge support upper bound for all vertices $u\in M_{i}$, it holds that $M_{i}.ub\_sup_{r}\geq u.ub\_sup_{r}$. Moreover, $u.ub\_sup_{r}$ is greater than the support of any edge in the $2r$-hop subgraph of $u$. Therefore, if $M_{i}.ub\_sup_{r}<k$ holds, the support of any edge connected to $u\in M_{i}$ is less than $k$ (for any user vertex $u\in M_{i}$), which indicates that $M_{i}$ can be safely pruned. ∎
Score Upper Bound Pruning for Synopsis Entries: Finally, we exploit the maximum upper bound, $M_{i}.ub\_score_{r}$, of user relationship scores and the score threshold $\sigma$ to discard synopsis entries.
Lemma 0.
(Synopsis Score Upper Bound Pruning) Given a synopsis entry $M_{i}$ and a user relationship score threshold $\sigma$, entry $M_{i}$ can be safely pruned, if it holds that $M_{i}.ub\_score_{r}<\sigma$.
Proof.
Since $M_{i}.ub\_score_{r}$ is the maximum score upper bound, it holds that $M_{i}.ub\_score_{r}\geq u.ub\_score_{r}$, where $u.ub\_score_{r}$ is the user relationship score upper bound of any user vertex $u\in M_{i}$. Moreover, it holds that $u.ub\_score_{r}\geq score_{u,u_{l}}(G_{t})$, where $score_{u,u_{l}}(G_{t})$ is the user relationship score between $u$ and its $2$-hop reachable user vertex $u_{l}$. By the inequality transition, we have $M_{i}.ub\_score_{r}\geq u.ub\_score_{r}\geq score_{u,u_{l}}(G_{t})$. If $M_{i}.ub\_score_{r}<\sigma$ holds, then we have $score_{u,u_{l}}(G_{t})<\sigma$ for all $u$ under $M_{i}$, which violates the score constraint. Therefore, in this case, we can safely prune entry $M_{i}$, which completes the proof. ∎
6.2. Snapshot CD-SBN Processing Algorithm
Algorithm 2 details our snapshot CD-SBN processing algorithm to obtain initial CD-SBN community answers, which consists of three major parts: i) the initialization of data structures and variables for preparing the synopsis traversal (lines 1-4); ii) the traversal of the synopsis, $Syn$, to obtain a candidate set $P$ (lines 5-18); and iii) the refinement of a snapshot CD-SBN result set $R$ (line 19).
Input: i) a streaming bipartite network $G_{t}$; ii) a tree synopsis $Syn$ over $G_{t}$, and; iii) a snapshot CD-SBN query with a set, $Q$, of query keywords, a query support threshold $k$, a query maximum radius $r$, and a threshold, $\sigma$, of user relationship score
Output: a set, $R$, of CD-SBN communities
1
// Initialization
2 hash all query keywords in $Q$ into a bit vector $Q.BV$;
3
4initialize a maximum heap $\mathcal{H}$ in the form of $(M,key)$;
5 insert $(root(Syn),+\infty)$ into heap $\mathcal{H}$;
6
[…]
8
// Synopsis Traversal
9 while $\mathcal{H}$ is not empty do
10       $(M,key)$ = $(\mathcal{H})$;
11       if $key<\sigma$ then // Lemma 6.3
12            terminate the loop;
13
14      if $M$ is a leaf node then
15             for each vertex $u_{i}\in M$ do
16                   if $2r$-hop subgraph $hop(u_{i},2r)$ cannot be pruned by Lemma 4.1, 4.2, 4.3, or 4.5 then
17                         obtain a maximal $(k,r,\sigma)$-bitruss $g\subseteq hop(u_{i},2r)$ satisfying the keyword constraint;
18                         if $g$ exists then
19                               add $g$ to $P$;
20
21
22
23
24      else // $M$ is a non-leaf node
25             for each entry $M_{i}\in M$ do
26                   if $M_{i}$ cannot be pruned by Lemma 6.1, 6.2, or 6.3 then
27                         insert entry $(M_{i},M_{i}.ub\_score_{r})$ into heap $\mathcal{H}$;
28
29
30
31
32
// Refinement
33 refine candidate communities in $P$ (removing the redundancy) and obtain the CD-SBN result set $R$;
34
35return $R$;
Algorithm 2 Snapshot CD-SBN Processing
Initialization: Given a query keyword set $Q$, the algorithm first hashes all query keywords in $Q$ into a query keyword bit vector $Q.BV$ (line 1). Then, we initialize a maximum heap $\mathcal{H}$ (for synopsis traversal), accepting heap entries in the form of $(M,key)$, where $M$ is a synopsis entry and $key$ is the maximum score upper bound, $M.ub\_score_{r}$, of all user vertices in $M$ (as mentioned in Section 5.3) (line 2). Intuitively, a heap entry with a large key (i.e., large maximum score upper bound) is more likely to contain communities with high user relationship scores. Next, we insert the root of the tree synopsis $Syn$, in the form $(root(Syn),+\infty)$, into the heap $\mathcal{H}$ (line 3). Moreover, we use an initial empty set $P$ to keep the candidate subgraphs and an empty set $R$ to store snapshot CD-SBN query answers (line 4).
Synopsis Traversal: Next, we utilize the maximum heap $\mathcal{H}$ to traverse the synopsis (lines 5-18). Each time we pop out a heap entry $(M,key)$ from heap $\mathcal{H}$ with the highest key, $key$ (i.e., the maximum score upper bound) (lines 5-6). If $key$ is less than the score threshold, $\sigma$, it indicates that all the remaining heap entries in $\mathcal{H}$ have a score less than $\sigma$, and they cannot be contained in our CD-SBN query results. Thus, we can safely terminate the synopsis traversal early (lines 7-8).
When $M$ is a leaf node, we consider each user vertex $u_{i}\in M$ and check whether or not we can find any candidate community centered on $u_{i}$ (lines 9-14). For each user $u_{i}\in M$, if the $2r$-hop subgraph $hop(u_{i},2r)$ cannot be pruned by our pruning strategies (i.e., keyword, support, and score upper bound pruning), then we will obtain a maximal $(k,r,\sigma)$-bitruss $g$ (satisfying the keyword constraint) within $hop(u_{i},2r)$ (lines 10-12). If such a subgraph $g$ exists, we will add $g$ to the candidate community set $P$ (lines 13-14).
On the other hand, when $M$ is a non-leaf node, we will check each entry $M_{i}$ of node $M$ (lines 15-16). If synopsis entries $M_{i}$ cannot be pruned using synopsis-level pruning strategies, then we will insert $(M_{i},M_{i}.ub\_score_{r})$ into the heap $\mathcal{H}$ for further checking (lines 17-18).
Refinement: Finally, we remove duplicate subgraphs from $P$ (due to overlap of $2r$-hop subgraphs), refine candidate communities in $P$ to obtain actual CD-SBN community answers in $R$, and return $R$ (lines 19-20).
Complexity Analysis: Let $PP^{(j)}$ be the pruning power (i.e., the percentage of node entries that can be pruned) at the $j$-th level of synopsis $Syn$, where $0\leq j\leq h$ (here $h$ is the height of the tree synopsis $Syn$). Denote $f$ as the average fanout of the synopsis nodes in $\mathcal{I}$. For the index traversal, the number of nodes that need to be accessed is given by $\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})$. Next, to obtain $g$ satisfying the constraints, we need to prune the item vertices without any query keyword. Therefore, it is necessary to recompute the support of the relevant edges and the user relationship score of any pair of adjacent user vertices. Let $\overline{n}$ be the number of pruned item vertices and $deg_{avg}$ be the average degree of the item vertices. The time cost of re-computation is $\overline{n}\cdot deg_{avg}^{2}$ (given in Section 5.2 and Section 5.4). Then, we can obtain the maximal $(k,r,\sigma)$-bitruss in $O(deg_{avg}^{2})$ with computed edge support, which is demonstrated in (Sarıyüce and Pinar, 2018). Thus, it takes $O(f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2})$ to refine candidate seed communities, where ${PP}^{(0)}$ is the pruning power over $r$-hop subgraphs in leaf nodes. Therefore, the total time complexity of Algorithm 2 is given by $O\left(\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})+f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2}\right)$.
6.3. Continuous CD-SBN Processing Algorithm
Algorithm 3 details our continuous CD-SBN query processing approach to monitoring answers. First, we initialize the CD-SBN result set $R_{t}$ at timestamp $t$ with that $R_{t-1}$ at timestamp $(t-1)$ (line 1). Next, we check the constraints of the existing communities to refine the previous result set (lines 2-6). Then, we obtain the potential communities from the $2r$-hop of user vertices near the insertion edge and remove the duplication (lines 7-10) Finally, we merge the candidate communities into the result set and return the result set at timestamp $t$ (lines 11-12).
Input: i) a streaming bipartite network $G_{t-1}$ with an update stream $S$; ii) a set, $R_{t-1}$, of community answers at timestamp $(t-1)$, and; iii) a continuous CD-SBN query with a set, $Q$, of query keywords, a support threshold $k$, a maximum radius $r$, and a threshold, $\sigma$, of user relationship score
Output: a set, $R_{t}$, of CD-SBN community answers at timestamp $t$
1
[…]
3
// Processing the Expired Item $p_{t-s}=(e_{u_{i},v_{a}},t-s)\in S$
4
5for each community answer $g\in R_{t}$ do
6       if $e_{u_{i},v_{a}}\in E(g)$ then
7             update the CD-SBN community $g$ with $g^{\prime}$, upon the weight update (or deletion) of edge $e_{u_{i},v_{a}}$;
8             if $g^{\prime}$ does not exist then
9                   remove $g$ from $R_{t}$;
10
11
12
// Processing the Insertion of a New Item $p_{t}=(e_{u_{i}^{\prime},v_{a}^{\prime}},t)\in S$
13 $\Delta R=\emptyset$;
14 for each user vertex $u_{j}\in hop(u_{i}^{\prime},2r)$ do
15       obtain a CD-SBN candidate community $g$ from $hop(u_{j},2r)$ (via pruning) and add it to $\Delta R$;
16
17remove duplicate communities $g$ from $\Delta R$ (overlapping with $R_{t}$) and refine candidate communities $g$ in $\Delta R$;
18 $R_{t}=R_{t}\cup\Delta R$;
19 return $R_{t}$;
Algorithm 3 Continuous CD-SBN Processing
Edge Expiration Processing: To deal with the expiration of an item $p_{t-s}$ ($=(e_{u_{i},v_{a}},t-s)$) in the sliding window $W_{t-1}$, the weight $w_{u_{i},v_{a}}$ of the edge $e_{u_{i},v_{a}}$ decreases (which might turn to the edge deletion). In this case, we will identify those community answers $g$ in the answer set $R_{t}$ that contain edge $e_{u_{i},v_{a}}$ (lines 2-3), and update the community $g$ with $g^{\prime}$ (upon weight/edge changes; line 4). If such a community $g^{\prime}$ does not exist (violating the CD-SBN predicates), we simply remove community answer $g$ from $R_{t}$ (lines 5-6).
Edge Insertion Processing: Next, for the insertion of a new item $p_{t}$ ($=(e_{u_{i}^{\prime},v_{a}^{\prime}},t)$), we aim to update $R_{t}$ with new CD-SBN community answers (containing the new/updated edge $e_{u_{i}^{\prime},v_{a}^{\prime}}$) in a set $\Delta R$ (lines 7-11). In particular, for each user vertex $u_{j}$ within the $2r$-hop away from $u_{i}^{\prime}$, we obtain a CD-SBN candidate community $g$ and add it to $\Delta R$ (if $g$ exists) (lines 8-9). Then, we will remove the redundancy of community candidates in $\Delta R$ (i.e., already in $R_{t}$) and refine the remaining candidate communities in $\Delta R$ (line 10). Finally, we add CD-SBN answers in $\Delta R$ to $R_{t}$, and return final CD-SBN answers in $R_{t}$ (lines 11-12).
Complexity Analysis: Let $deg_{avg}$ be the average degree of the item vertices. For edge insertion and expiration, the time cost of user relationship score re-computation is $O(deg_{avg})$ and of edge support update is $O(deg_{avg}^{2})$. Since obtaining the maximal $(k,r,\sigma)$-bitruss with computed edge supports and user relationship scores (Sarıyüce and Pinar, 2018) requires $O(deg_{avg}^{2})$, it takes $O(|R_{t-1}|deg_{avg}^{2})$ to maintain the result set at $t-1$ and $O(deg_{avg}^{3})$ to detect potential communities. Therefore, the total time complexity of Algorithm 3 at each timestamp $t$ is given by $O\left(|R_{t-1}|deg_{avg}^{2}+deg_{avg}^{3}\right)$.
7. Experimental Evaluation
7.1. Experimental Settings
We conduct a set of experiments to examine the performance of our snapshot and continuous CD-SBN query processing algorithms over different real-world bipartite graphs. Our source code is available on GitHub [^1] https://github.com/L1ANLab/CD-SBN.
Table 2. Statistics of graph data sets.
Data Set Type $|U|$ $|V|$ $|E|$ $|\Sigma|$
AmazonW (AM) Rating 26,112 800 111,265 N/A
BibSonomy (BS) Publication 5,795 767,448 2,555,080 204,674
CiaoDVD (CM) Rating 17,616 16,121 295,958 N/A
CiteULike (CU) Publication 22,716 731,770 2,411,819 153,278
Movielens (ML) Movie 4,010 7,602 95,580 16,529
Escorts (SX) Rating 10,106 6,624 50,632 N/A
TripAdvisor (TA) Rating 145,317 1,760 703,171 N/A
UCForum (UF) Interaction 899 522 33720 N/A
ViSualizeUs (VU) Picture 17,110 495,402 2,298,816 82,036
dBkU, dBkL, dBkP Synthetic 25,000 25,000 152,175 N/A
dPkU, dPkL, dPkP Synthetic 25,000 25,000 152,922 N/A
Bipartite Graph Data Sets: We evaluate our proposed CD-SBN algorithms on nine real-world and six synthetic graphs. In particular, as depicted in Table 2, we provide statistics of real-world bipartite graphs from the KONECT [^2] http://konect.cc/ project (i.e., Koblenz Network Collection), where $|\Sigma|$ refers to the domain size of keywords in item vertices of the original graphs (“N/A” for $|\Sigma|$ means that graph data have no keywords).
To generate synthetic bipartite graphs, we first randomly produce degrees, $deg(u)$, of user vertices $u$ following PowerLaw or Beta distribution. Then, for each user vertex $u$, we connect it to $deg(u)$ random item vertices $v$. Next, to simulate the real-world frequency of user-item interaction (i.e., edge weights $w_{u,v}$), we assign each edge with weight ranging from $\left[min\_w,max\_w\right]$ following the Gaussian distribution, and associate them with edges, where the mean and standard deviation of the Gaussian distribution are $(1.5,0.25)$, $(2,0.5)$, and $(2.5,0.75)$ for $\left[1,2\right]$, $\left[1,3\right]$, and $\left[1,4\right]$, respectively. Finally, for each item vertex $v$, we generate its keyword set $v.K$, which contains integers following Log-Normal, Pareto, or Uniform distribution. For real-world graphs without keywords (e.g., AM, CM, SX, UF, and TA), we produce keyword sets of their item vertices following the Log-Normal distribution.
For CD-SBN predicates, we will randomly select $|Q|$ keywords from a keyword domain $\Sigma$, following the keyword distribution in the data graph $G$, and form a set, $Q$, of query keywords. Other parameter settings are provided in Table 3.
Competitors: Since no prior works studied the CD-SBN problem under the same semantics as $(k,r,\sigma)$-bitruss communities with specific keywords, we compare our CD-SBN algorithm with a baseline method, named Bitrussness-Based Decomposition (BBD). Specifically, BBD first calculates the trussness of each edge, which is the maximum $k$ of all the $k$-bitruss containing the edge and builds a synopsis based on the trussness. Then, upon edge updates at each timestamp, BBD filters out user vertices connected to edges with trussness less than $k$ and computes the $(k,r,\sigma)$-bitruss community centered at each remaining vertex (containing query keywords).
Table 3. Parameter settings.
Parameters Values
sliding window size $s$ 200, 300 500, 800, 1000
support, $k$, of bitruss structure 3, 4, 5
radius $r$ 1, 2, 3
relationship score threshold $\sigma$ 1, 2, 3, 4, 5
size, $|Q|$, of query keyword set $Q$ 2, 3, 5, 8, 10
keyword domain size $|\Sigma|$ 100, 200, 500, 800, 1000
size, $|v_{i}.K|$, of keywords per item vertex 1, 2, 3, 4, 5
the number, $|U(G)|$, of user vertices 10K, 15K, 25K, 50K, 100K
the number, $|L(G)|$, of item vertices 10K, 15K, 25K, 50K, 100K
the edge weight range $\left[min\_w,max\_w\right]$ [1,2], [1,3], [1,4]
Measures: We evaluate our CD-SBN performance, in terms of the wall clock time, which is the time cost to traverse the index and retrieve communities for snapshot CD-SBN, or that to maintain the CD-SBN results at each timestamp for continuous CD-SBN problem. The wall clock time is an average time cost over 10 runs with a set of 10 different CD-SBN predicates (e.g., query keywords).
(a) real-world graphs
(b) synthetic graphs
Figure 3. CD-SBN performance on data sets.
Parameter Settings: Table 3 depicts the parameter settings of our subsequent experiments, where default values are in bold. Each time we vary one parameter while setting other parameters to their default values. Moreover, since our continuous CD-SBN processing algorithm has no data dependencies during the search of communities, we use multithreaded optimization to speed up the computation, where the number of threads is set to $20$ by default. All the experiments are executed on a machine with an AMD Ryzen Threadripper 3990X 64-Core processor, Ubuntu 22.04 OS, 192GB memory, and a 600GB disk. All algorithms were implemented in C++ (C++17 standard) and compiled with the g++ 11.4.0 compiler.
Research Questions: We design experiments for our CD-SBN algorithms to answer the following research questions (RQs):
RQ1 (Efficiency): Can our proposed algorithms efficiently process continuous CD-SBN problem?
RQ2 (Effectiveness): Can our proposed pruning strategies effectively filter out false alarms of candidate communities during the CD-SBN processing?
RQ3 (Utility): Do the resulting CD-SBN communities have practical significance for real-world applications?
(a) sliding window size $s$
(b) bitruss support threshold $k$
(c) radius $r$
(d) relationship score threshold $\sigma$
(e) query keyword set size $|Q|$
(f) keyword domain size $|\Sigma|$
(g) # of keywords/vertex $|v_{i}.K|$
(h) edge weight range
(i) lower layer size $|L(G)|$
(j) upper layer size $|U(G)|$
Figure 4. Robustness evaluation of CD-SBN.
7.2. CD-SBN Performance Evaluation
Continuous CD-SBN Efficiency (RQ1): Figure 3 compares the wall clock time of our continuous CD-SBN processing algorithm with that of BBD over real-world and synthetic bipartite graphs, where all parameters are set to default values in Table 3. The experiment results indicate that our continuous CD-SBN algorithm outperforms BBD by more than two orders of magnitude, confirming the efficiency of our CD-SBN approach on real/synthetic graphs.
We also test the robustness of our CD-SBN approach with different parameters (e.g., $s$, $k$, $r$, $\sigma$, and $|Q|$) on synthetic graphs with different structural parameters (e.g., $|\Sigma|$, $|v_{i}.K|$, $[min\_w,max\_w]$, $|L(G)|$, and $|U(G)|$).
Effect of Sliding Window Size $s$: Figure 4(a) presents the performance of our continuous CD-SBN approach, by varying the size $s$ of sliding window $W_{t}$ from $50$ to $200$, where other parameters are by default. From the experimental results, with the increasing $s$, more edges are involved or have higher edge weights, which result in more potential candidate communities. Therefore, larger window size $s$ incurs higher time cost. Nonetheless, the time cost of our CD-SBN approach remains low (i.e.,0.047 $\sim$ 0.726 $sec$) for different $s$ values.
Effect of Bitruss Support Threshold $k$: Figure 4(b) illustrates the continuous CD-SBN performance, where the bitruss support parameter $k$ varies from $3$ to $5$, and other parameters are set to default values. Since a larger support threshold $k$ leads to higher pruning power, our CD-SBN approach can obtain fewer candidate communities, which in turn incurs lower time costs. The query cost remains low for different $k$ values (i.e., 0.053 $\sim$ 0.277 $sec$).
Effect of Radius $r$: Figure 4(c) shows the experimental results of our CD-SBN approach, by varying radius, $r$, the query community from $1$ to $3$, where default values are used for other parameters. When radius $r$ becomes larger, the scale of $2r$-hop subgraphs will increase, which results in higher time costs to obtain maximal $(k,r,\sigma)$-bitruss. Nevertheless, for different $r$ values, the wall clock time remains small (i.e., around 0.033 $\sim$ 0.375 $sec$).
Effect of Relationship Score Threshold $\sigma$: Figure 4(d) varies the relationship score threshold, $\sigma$, from $1$ to $5$, where other parameters are set by default values. Since higher $\sigma$ leads to higher pruning power and cohesiveness of candidate communities, the number of candidate communities decreases for larger $\sigma$. Thus, the CD-SBN time cost decreases for larger $\sigma$ values. Overall, for different $\sigma$ values, the time cost remains low (i.e., 0.051$\sim$1.989 $sec$).
Effect of Query Keyword Set Size $|Q|$: Figure 4(e) evaluates the CD-SBN performance, with different sizes, $|Q|$, of the query keyword set $Q$ from $2$ to $10$, where we use default values for other parameters. Intuitively, a larger query keyword set $Q$ allows more item vertices to be included (with lower pruning power by the keyword pruning), and in turn more candidate communities will be retrieved for filtering/refinement. Thus, from the figure, when $|Q|$ becomes larger, the query cost also increases. Nonetheless, the wall clock time remains low (i.e., 0.007 $\sim$ 0.893 $sec$) for different $|Q|$ values.
Effect of Keyword Domain Size $|\Sigma|$: Figure 4(f) shows the wall clock time of our CD-SBN approach, where the keyword domain size, $|\Sigma|$, varies from $100$ to $1,000$ and other parameters are set to their default values. In the figure, we can see that, when $|\Sigma|$ becomes larger, our CD-SBN approach requires less time to retrieve CD-SBN communities, which indicates the effectiveness of our keyword pruning for large keyword domain size. Overall, with different $|\Sigma|$ values, the time cost remains low (i.e., 1.974 $\sim$ 0.009 $sec$).
Effect of # of Keywords Per Item Vertex $|v_{i}.K|$: Figure 4(g) illustrates the CD-SBN performance for different numbers, $|v_{i}.K|$, of keywords per item vertex from $1$ to $5$, where we set other parameters to default values. Intuitively, more keywords in item vertices $v_{i}$ will increase the chance to match with query keyword set $Q$, which results in more or larger candidate communities to filter and refine. Therefore, with the increase of $|v_{i}.K|$, the time cost of our CD-SBN approach also increases, but, nevertheless, remains low (i.e., 0.014$\sim$0.821 $sec$).
Effect of Edge Weight Range $\left[min\_w,max\_w\right]$: Figure 4(h) examines the CD-SBN performance with different edge weight ranges, $\left[min\_w,max\_w\right]$ (i.e., $[1,2]$, $[1,3]$, and $[1,4]$), where default values are used for other parameters. When the range of edge weights expands, more/larger candidate communities will be retrieved and refined. Thus, the query cost increases as the edge weight interval $\left[min\_w,max\_w\right]$ becomes wider. Nonetheless, the wall clock time remains low (i.e., 0.0543$\sim$2.761 $sec$).
Effect of Lower Layer Size $|L(G)|$: Figure 4(i) tests the scalability of our proposed CD-SBN approach by varying the lower layer size, $|L(G)|$, from $10K$ to $100K$, where default values are used for other parameters. From the figure, we can see that the query cost decreases as $|L(G)|$ increases. This is because, with a constant average degree of each user vertex, larger $|L(G)|$ leads to a more sparse bipartite network, which incurs fewer candidate communities and lower query costs. The time cost remains low (i.e., 0.031$\sim$0.407 $sec$).
Effect of Upper Layer Size $|U(G)|$: Figure 4(j) evaluates the scalability of our CD-SBN approach with different upper layer sizes, $|U(G)|$, from $10K$ to $100K$, where other parameters are set by default. In the figure, with the increase of $|U(G)|$, the CD-SBN time cost increases due to more retrieved candidate communities. Nonetheless, the time cost remains low (i.e., 0.001 $\sim$ 1.815 $sec$), which confirms the good scalability of our CD-SBN approach for large $|U(G)|$.
Ablation Study on the Snapshot CD-SBN Pruning Strategies (RQ2): To test the effectiveness of our proposed pruning methods, we conduct an ablation study of our snapshot CD-SBN pruning strategies over real-world/synthetic graphs, where all parameters are set to default values. We execute the snapshot CD-SBN processing algorithm by adding one pruning strategy at a time, which forms 3 experimental groups: (1) support pruning only, (2) support + score pruning, and (3) support + score + keyword pruning. Figure 5 reports the number of pruned candidate communities by using different pruning method groups. In this figure, we can see that for all data sets, the number of pruned communities increases by about an order of magnitude, as we apply more pruning strategies, which indicates the effectiveness of our proposed pruning strategies.
Figure 5. Ablation study of the pruning power for the snapshot CD-SBN problem.
Figure 6. Case study of $(k,r,\sigma)$-bitruss vs. biclique (Lin, 2018).
CD-SBN Community Case Study (RQ3): To evaluate the utility of our CD-SBN results, we carry out a case study to compare our proposed $(k,r,\sigma)$-bitruss pattern with the biclique (Lin, 2018) (which is a variant of bipartite community structure) over BibSonomy (BS) graph data. The entire graph shown in Figure 6 is a $(4,2,3)$-bitruss containing 3 users (in blue) and 9 items (in blue or red), whereas the 3 users, 4 items (in blue) form a $3\times 4$ biclique. We can see that the $3\times 4$ biclique shares the same 3 users as that in $(4,2,3)$-bitruss, but the $3\times 4$ biclique ignores some items that have edges connected to some users with large weights (e.g., 6). In contrast, our $(k,r,\sigma)$-bitruss captures these items, and can recommend them to users who are not connected to them (e.g., recommending items $v\_449272$, $v\_449273$, $v\_449277$, $v\_449275$ and $449276$ to user $u\_4954$). This confirms the utility of our proposed CD-SBN community semantics, which can recommend items to more potentially interested users.
8. Related Work
Community Search/Detection Over Unipartite Graphs: As fundamental operations in social network analysis, the community search (CS) and community detection (CD) have been extensively studied for the past few decades (Al-Baghdadi and Lian, 2020; Huang and Lakshmanan, 2017; Sozio and Gionis, 2010; Cui et al., 2014, 2013; Newman and Girvan, 2004; Fortunato, 2010; Xu et al., 2012; Conte et al., 2018; Veldt et al., 2018), which obtain communities that satisfy different query parameters (e.g., keyword relevance and structural cohesiveness). In unipartite graphs, such as social networks, different community semantics have been proposed, including minimum degree (Cui et al., 2014), $k$-core (Sozio and Gionis, 2010), $k$-clique (Cui et al., 2013), and $(k,d)$-truss (Huang and Lakshmanan, 2017; Al-Baghdadi and Lian, 2020). Previous works on community detection usually retrieved all communities considering link information only (Newman and Girvan, 2004; Fortunato, 2010). Some recent works used clustering techniques to detect communities (Xu et al., 2012; Conte et al., 2018; Veldt et al., 2018). In contrast, our work is conducted on bipartite graphs with two distinct types of nodes (instead of unipartite graphs with the same type of nodes). Moreover, our proposed $(k,r,\sigma)$-bitruss in bipartite graphs considers not only the high connectivity of community structure but also edge weight and query keywords, which is more challenging. Thus, previous work on CS/CD over unipartite graphs cannot be used for streaming bipartite graphs (i.e., focus of our CD-SBN problem).
Dynamic Community Search/Detection Over Unipartite Graphs: In recent years, due to the importance of mining hidden trends or patterns in streaming data, community search/detection over large-scale dynamic graphs has become an increasingly important research problem. Some existing studies aim to find potential communities over dynamic social networks with different specified community structures to gain valuable insights, e.g., over unipartite dynamic graphs, (Xu et al., 2022a) searches for and maintains a Triangle-connected $k$-Truss Community where edges should be contained in adjacent triangles and (Tang et al., 2024, 2022) work on the $k$-core with edge weights higher than $\theta$ (i.e., $(\theta,k)$-core). Unlike the existing tasks, our proposed CD-SBN problem focuses on detecting communities over streaming bipartite graphs, requiring more complicated calculations and greater computational overhead.
Community Search/Detection Over Bipartite Graphs: Several prior works on the CS/CD problem over bipartite graphs considered different community semantics, including $(\alpha,\beta)$-core community (Wang et al., 2021), size-bounded $(\alpha,\beta)$-community (Zhang et al., 2024), personalized maximum biclique community (Wang et al., 2022), and $k$-wing community (Abidi et al., 2023a). However, since these works are focused on searching potential communities over static bipartite graphs, we cannot directly apply previous techniques to solve our CD-SBN problem with dynamically changing bipartite graphs. This is because our proposed CD-SBN problem focuses on a different community semantic, i.e., $(k,r,\sigma)$-bitruss, which includes the constraints of structural cohesiveness, keyword matching, and user relationship scores. On the other hand, few prior works (Wang et al., 2023; Abidi et al., 2023b) consider detecting or searching communities on dynamic streaming bipartite graphs. Although all of them construct an index that can be dynamically updated, for each query, these methods still need to re-traverse the entire index. In contrast, our CD-SBN approach only needs to incrementally maintain the result set upon edge updates, which significantly reduces the query cost.
9. Conclusions
In this paper, we propose a novel CD-SBN problem to detect those communities with user-specified query keywords and high structural cohesiveness over streaming bipartite graphs in snapshot and continuous scenarios. To efficiently tackle the CD-SBN problem, we provide effective pruning strategies to rule out false alarms of candidate communities and design a hierarchical synopsis to facilitate the CD-SBN processing. We also develop efficient algorithms to enable snapshot/continuous CD-SBN answer retrieval by quickly accessing the synopsis and applying our proposed pruning methods. Comprehensive experiments confirmed the effectiveness and efficiency of our proposed CD-SBN approaches over real/synthetic streaming bipartite networks upon graph updates.